OpenStudy (anonymous):

Will give medal and fan.... Shannon finances $189,000 with a 20/7 balloon mortgage at 6.25%. How much will she pay for principal and interest over the life of the loan? $261,951.94 $264,100.53 $266,751.42 $497,322.00 I keep getting 190,383.66... I am not sure what I am doing wrong.

OpenStudy (anonymous):

I am using the formula P[((i(1+i)^((n)(t)))/((1+i)^((n)(t))-1)].... where P= 189,000 i= .0625/12 n=12 t=20

OpenStudy (anonymous):

@amistre64 can you please help me? (:

OpenStudy (amistre64):

add up all the payments right?

OpenStudy (amistre64):

7 years of payments, plus the remaining amount after 7 years for a balloon payment

OpenStudy (amistre64):

you have a 20 year fixed loan, for 7 years ... add up the payments after the 7 years is up, there is an amount left on the loan that gets paid off ..

OpenStudy (amistre64):

7*12 payments of 1381.45 each + a remaining balance of 147,292

OpenStudy (anonymous):

So, for 7 years there would be no interested added? Then after those 7 years are up you would then have a fixed interest rate for 20 years... and you solve for the 7 years then you solve for the 20 years separately then add the total of those two together?

OpenStudy (amistre64):

interest is always applied to a house loan unless told other wise the total number of payments is the total cost of the loan .. principal + interest

OpenStudy (amistre64):

if you know the monthly payments on a 20 yr fixed loan ... then after 7 years youve made how much in payments? the loan is to be paid in full after 7 years, so we need to determine the remaining balance of the loan after 7 years to make a 1 time payment of ....

OpenStudy (amistre64):

\[(7*12)*P~+~Bk^{7*12}-P\frac{1-k^{7*12}}{1-k}\]

OpenStudy (amistre64):

factoring out a P if you want to simplify it and such ....\[Bk^{84}+P~(84-\frac{1-k^{84}}{1-k})\]

OpenStudy (amistre64):

im getting 263,334 but am prolly being too precise so id go with the 264 option

OpenStudy (anonymous):

Yes that is what I got as well... was 263,334... it was taking me a while while to respond because I was checking my work xp

OpenStudy (anonymous):

Thank you so much for your help (:

OpenStudy (amistre64):

youre welcome